# The Developments and Challenges towards Dexterous and Embodied Robotic Manipulation: A Survey

### v1

@@ Abstract
Index Terms:
Dexterous Manipulation, Multi-fingered Hands, AI-Enabled Robotics, Data Collection, Imitation Learning, Reinforcement Learning.
@@ I Introduction
However, even today, robotic manipulation is still an unresolved and open problem. The performance of robotic manipulation is still far from humans’ expectation, especially when it comes to hand’s dexterous manipulation[12][13]. Therefore, robotic manipulation is still facing numerous challenges and is a hot topic at the cutting edge of the robotics field.
@@ II The Historical Stages of Robotic Manipulation
In recent years, a consensus has been reached that the disembodied intelligence without a physical body will eventually reach a ceiling, and the embodied intelligence with physical entities (e.g., robots) to interact with the world is the only way to achieve general AI[21]. The development of robotic manipulation can be divided into three stages as shown in Fig. 2.
@@
The second stage is the Closed-Loop Control Stage based on visual servo. The visual feedback is introduced into the control loop through the Eye-in-Hand or Eye-to-Hand camera to achieve closed-loop control based on feature tracking. This enables robotic manipulation to have a certain degree of adaptability to changes in tasks and environments. Typical achievements such as the ABB YuMi dual-arm collaborative robot [23], which can complete component assembly tasks in a semi-structured environment. However, the control of the robot at this stage still relies on accurate modeling of the external environment or the operated workpiece. For example, in a model-based grasping task, force closure detection is a necessary condition for determining whether a stable grasp can be formed. However, the inaccuracy of the model will greatly reduce the manipulation performance, while the unstructured environment is always filled with a large number of objects and disturbances that are difficult to be model in advance.
With the development of AI, robotic manipulation has gradually entered the Embodied Intelligent Manipulation Stage that emphasizes the end-to-end "perception-decision-execution" closed loop [24]. Embodied intelligent manipulation enables robots to enhance their understanding of the surrounding environment based on multi-modal sensor information such as vision, force, and tactile, and is expected to enable robots to adapt to dynamic unstructured environments. It is currently the most promising technical route to give robots the human-like dexterous manipulation abilities.
@@ III The Challenges in Dexterous Manipulation: from Gripper to Muti-Fingered Hand
In the embodied intelligent manipulation stage, the complexity of interacting with the physical world is the main source of the severe challenges faced by robotic manipulation. As robots move beyond structured factory environments, the objects manipulated by robots have also changed from traditional rigid workpieces to complex objects that are difficult to be sense or model. As shown in Fig. 3, it is challenging for robots to manipulate linear objects (e.g., cables)[25], transparent objects (e.g., glasses)[26], soft objects (e.g., cloth)[27] [28], and deformable objects (e.g., human tissues in medical operations)[29]. When robots interact with the physical world, except for the more and more complex manipulated objects, the types of robotic manipulation have also become more diverse. In addition to the basic pick-and-place manipulation, a lot of robotic manipulation types are required for interaction, such as point-touch sliding manipulation that relies on single-point contact to achieve complex object movement[30], push-pull manipulation (e.g. turning an abacus and opening/closing switches)[31], in-hand manipulation that only relies on one hand to achieve object rotation[32], and dexterous manipulation in various scenarios (e.g., opening/closing valves, grasping objects, and using tools)[33]. Complex objects and diverse manipulations make robotic manipulation face challenges in difficult modeling, high uncertainty, large disturbances, and limited sense capabilities, which greatly affect the further expansion and application of robot technology.
@@
How to solve these problems is the frontier and hot topic in current research. However, current research on robotic manipulation mainly focuses on two-fingered grippers. For example, the robotic large language model (LLM) RT-X series proposed by Google [34] and VoxPoser [35] which combines LLM and vision-language model (VLM) for robotic manipulation. Although a simple two-fingered gripper can finish many manipulation tasks, this type of end-effector has two disadvantages which seriously restrict its development in robotic manipulation. First, the two-fingered gripper can only provide fewer contact points during the interaction and cannot provide stable and reliable gripping under certain tasks. Second, the two-fingered gripper itself has fewer degrees of freedom and cannot complete complex manipulation tasks, such as in-hand rotating and using tools. Highly anthropomorphic multi-fingered dexterous hands can make up the deficiency of two-fingered grippers.
@@
Compared with the simple two-fingered gripper, the learning of manipulation skills for multi-fingered dexterous hands is particularly difficult. First, the multi-fingered dexterous hand has a higher degree of freedom (DoF) and a more variable interaction space, which greatly increases the difficulty of searching in high-dimensional space for algorithms such as RL. Second, in the process of performing manipulation tasks, the multi-fingered dexterous hand has multiple contact points with objects, leading to more complex and diverse mechanical properties. This further makes the learning of skills for multi-fingered dexterous hands more difficult.
@@ IV-A Data Generation Based on Simulation Platform
Faced with complex, diverse, and highly uncertain application scenarios, traditional learning methods that rely on "robot-environment" interaction to obtain intelligence often require millions of iterations to learn useful skills, which is generally inefficient. Therefore, the most common approach is to generate a large amount of data through simulation platforms to improve efficiency. Data generation based on simulation platforms has some unique advantages. First, generating data on the simulation platform can avoid the high cost and potential danger of repeated experiments in the real world. Second, the simulation platform can efficiently, cost-effectively and repeatably generate a large amount of pedigree data with comparative significance. This can enhance the diversity of datasets and improve the quality of generated data. Finally, the simulation platform can accelerate the speed of data collection that may take longer in the real physical world by changing the flow rate of time in simulation environment. At present, there have developed some simulation datasets based on physics engines such as Genesis[50], Isaac Sim, PyBullet[51], and MoJoCo[52]. For example, the GraspM3[37] simulation dataset that provides millions of grasping trajectories covering 8,152 objects and a generative robot agent RoboGen[38] that can generate data infinitely. Wang et al.[39] generated a large-scale simulation dataset DexGraspNet for Shadow Hand, which contains 1.32 million grasping data of 5,355 objects in 133 categories. Shanghai Artificial Intelligence Laboratory[40] released a city-level embodied intelligent simulation platform GRPtopia, which can provide 89 functional scenarios and 100,000 high-quality interactive data for robot training.
@@ IV-B Data Collection from Human Demonstration
The ultimate goal of robotic manipulation is to endow robots with human-like dexterous manipulation capabilities. Therefore, learning manipulation skills directly from human demonstration actions has become an attractive solution. Although the scale of data collection based on human motion demonstrations cannot compare with the data generated by simulation, it significantly reduces the difficulty compared to repeated trials with real robots, and can expand the scope of data collection without requiring robot hardware. At the same time, compared with the data generated by simulation, the human demonstration data is real interaction data produced in the physical environment, which can greatly reduce the Sim2Real gap. For instance, Bahl et al.[41][42] proposed to extract human motions and key interaction elements from billions of video data that are already widely available in Internet to achieve manipulation data collection based on human demonstration. Liu et al.[43][44] also developed a vision-tactile fusion human motion capture system, constructing the VTDexManip dataset. These works enhance robots’ manipulation capabilities in real physical environments.
@@ V The Learning Framework for Dexterous Manipulation Skills
BC [65] Rely on a single RGB camera solely to capture human demonstrations via teleoperation of a robotic hand.
BC [66] Utilize implicit function to generate continuous grasping motions for dexterous hands from human demonstrations.
BC [67] Combine behavior classification and geometric augmentation to enable effective behavioral cloning from mixed-quality demonstration datasets.
IRL [68] Utilize a pre-trained visual dynamics model to learn cost functions when given only visual human demonstrations. The learned cost functions are then used to reproduce the demonstrated behavior via visual model predictive control.
IRL [69] Combine random sample generation and reward normalization to overcome demonstration-dependent reward bias.
GAIL [70] Recover the expert’s cost function with IRL, then extract a policy from that cost function with RL.
DIH-Tele [71] Tactile information is utilized for IL of a multi-fingered dexterous hand.
@@
DAPG [75] First work and the introduction of pre-trained model.
@@
— — [77] Use only visual tactile sensing inputs for RL. Policy that obtained by RL can be generalized to unseen objects.
@@
— — [81] DRL is augmented with a small number of human demonstrations.
DexH2R [82] Human hand motion retargeting and a task-oriented residual action policy are combined.
@@
— — [84] A Bayesian-inspired lifelong RL framework integrates language embeddings to enable robots to continuously accumulate and reuse knowledge from sequential tasks, advancing generalizable robotic intelligence.
@@ V-A Imitation Learning
The second subcategory of IL employs deep learning to train policy networks that mimic human expert decisions. Unlike RL, this approach does not rely on environmental interaction or rewards but learns from pre-collected datasets. Common methods include Behavior Cloning (BC)[65, 66, 67, 71], Inverse reinforcement learning (IRL)[68, 69], and Generative Adversarial Imitation Learning (GAIL)[70]. A key limitation is that these methods only learn static behaviors from the demonstration data and cannot surpass human performance as RL might. Additionally, their performance degrades significantly when encountering states outside the training distribution, highlighting robustness issues.
@@ V-B Reinforcement Learning
RL has become the dominant approach for robotic dexterous manipulation skill learning, enabling agents to interact with their environment and refine policies through reward feedback. Recent researches have developed specialized simulation environments (e.g., Bi-DexHands for bi-manual manipulation)[72], and RL methods for robotic control. Compared with traditional feedback control methods, RL has successfully tackled challenging manipulation skills[73, 85, 74]. However, applying RL to multi-fingered dexterous hands presents unique challenges due to the high degrees of freedom and complex contact dynamics[86]. Pure RL that only relies on agent-environment often suffers from sparse rewards and low sample efficiency, typically requiring millions of iterations to learn useful behaviors. The exploration process can also lead to incoherent or unsafe actions, making it difficult to achieve desired learning outcomes. To enhance learning efficiency, recent studies have utilized features extracted by pre-trained models to guide downstream robot tasks[75, 76]. The use of existing pre-trained models, combined with vast internet data, enables rapid and efficient learning of new tasks. In robotic RL, this pre-trained paradigm offers a key advantage: robots can acquire new skills with minimal human intervention by leveraging publicly available data and pre-trained models. Same as IL, tactile information can be quite useful. Su et al. employed RL to train a robot to pivot an object to a target orientation using only proprioceptive and tactile information[77]. Liu et al. proposed a novel framework called TactileAIRL for skill learning in robotic dexterous manipulation. Model-based techniques and intrinsic curiosity are integrated into the RL process, the method leverages vision-based tactile sensing to extract meaningful contact features. This design makes TactileAIRL scalable to many manipulation tasks learning involving tactile feedback[78]. Zhou et al. presented an RL framework called T-TD3 for stable grasping of deformable objects using tactile prior information[79]. Aslam et al. presented a robot called DartBot that integrates tactile exploration and RL to achieve robust throwing skills in object transfer through throwing task[80].

### v2

@@ I Introduction
However, even today, robotic manipulation is still an unresolved and open problem. The performance of robotic manipulation is still far from humans’ expectation, especially when it comes to hand’s dexterous manipulation[12][13]. On one hand, the dexterous hands have high degrees of freedom (DoFs), making it difficult to achieve precise control through models. On the other hand, the learning-based methods need a large amount of high-quality datasets for training, especially the human demonstration data which can significantly accelerate the training process. Therefore, robotic manipulation is still facing numerous challenges and is a hot topic at the cutting edge of the robotics field.
@@ II The Historical Stages of Robotic Manipulation
In recent years, a consensus has been reached that the disembodied intelligence without a physical body will eventually reach a ceiling, and the embodied intelligence with physical entities (e.g., robots) to interact with the world is the only way to achieve general AI[21]. Meanwhile, the robotic manipulation has entered the stage of embodied intelligence. This evolution stems not only from advances in AI and robotic hardware, but also from growing performance expectations for robotic manipulation itself. To better understand its characteristics and pathway to embodied intelligence, this section examines the developmental history of robotic manipulation. The development can be divided into three stages as shown in Fig. 2.
@@
The second stage is the Closed-Loop Control Stage based on visual servo. The visual feedback is introduced into the control loop through the Eye-in-Hand or Eye-to-Hand camera to achieve closed-loop control based on feature tracking. This enables robotic manipulation to have a certain degree of adaptability to changes in tasks and environments. Typical achievements such as the ABB YuMi dual-arm collaborative robot [23], which can complete component assembly tasks in a semi-structured environment. However, the control of the robot at this stage still relies on accurate modeling of the external environment or the operated workpiece. For example, in a model-based grasping task, force closure detection is a necessary condition for determining whether a stable grasp can be formed. However, the inaccuracy of the model will greatly reduce the manipulation performance, while the unstructured environment is always filled with a large number of objects and disturbances that are difficult to be modeled in advance.
With the development of AI, robotic manipulation has gradually entered the Embodied Intelligent Manipulation Stage that emphasizes the end-to-end "perception-decision-execution" closed loop [24]. This paradigm represents a fundamental shift from previous stages. On the perception front, robots no longer rely solely on visual feedback but instead leverage multi-modal sensor fusion, integrating vision, force, and tactile sensing. Force and tactile feedback provide critical information about physical interactions, enabling manipulation in unstructured environments where visual feedback alone are insufficient. On the decision front, robotic manipulation is gradually breaking free from the constraints of precise environmental modeling. Robot agents can now autonomously decide their next actions and trajectories based on real-time multi-modal perceptual feedback, even in previously unknown environments. On the execution front, robots are increasingly equipped with multi-fingered dexterous hands as end-effectors, which are capable of acquiring complex manipulation skills and achieving far more flexible and adaptive interactions with objects. Embodied intelligent manipulation enables robots to enhance their understanding of the surrounding environment based on multi-modal sensor information such as vision, force, and tactile, and is expected to enable robots to adapt to dynamic unstructured environments. It is currently the most promising technical route to give robots the human-like dexterous manipulation abilities.
@@ III The Challenges in Dexterous Manipulation: from Gripper to Muti-Fingered Hand
In the embodied intelligent manipulation stage, the complexity of interacting with the physical world is the main source of the severe challenges faced by robotic manipulation. As robots move beyond structured factory environments, the objects manipulated by robots have also changed from traditional rigid workpieces to complex objects that are difficult to be sensed or modeled. As shown in Fig. 3, it is challenging for robots to manipulate linear objects (e.g., cables)[25], transparent objects (e.g., glasses)[26], soft objects (e.g., cloth)[27] [28], and deformable objects (e.g., human tissues in medical operations)[29]. When robots interact with the physical world, except for the more and more complex manipulated objects, the types of robotic manipulation have also become more diverse. In addition to the basic pick-and-place manipulation, a lot of robotic manipulation types are required for interaction, such as point-touch sliding manipulation that relies on single-point contact to achieve complex object movement[30], push-pull manipulation (e.g. turning an abacus and opening/closing switches)[31], in-hand manipulation that only relies on one hand to achieve object rotation[32], and dexterous manipulation in various scenarios (e.g., opening/closing valves, grasping objects, and using tools)[33]. Complex objects and diverse manipulations make robotic manipulation face challenges in difficult modeling, high uncertainty, large disturbances, and limited sense capabilities, which greatly affect the further expansion and application of robot technology.
@@
How to solve these problems is the frontier and hot topic in current research. However, current research on robotic manipulation mainly focuses on two-fingered grippers. For example, the robotic large language model (LLM) RT-X series proposed by Google [34] and VoxPoser [35] which combines LLM and vision-language model (VLM) for robotic manipulation. Although a simple two-fingered gripper can finish many manipulation tasks, this type of end-effector has two disadvantages which seriously restrict its development in robotic manipulation. First, the two-fingered gripper can only provide fewer contact points during the interaction and cannot provide stable and reliable gripping under certain tasks. Second, the two-fingered gripper itself has fewer degrees of freedom and cannot complete complex manipulation tasks, such as in-hand rotating and using tools.
Highly anthropomorphic multi-fingered dexterous hands can make up the deficiency of two-fingered grippers. On the one hand, it has more DoFs, contact points, and a larger operating space, which enhance adaptability and stability in uncertain environments and when handling objects difficult to model. On the other hand, dexterous hands support a wider range of skills instead of only opening and closing as grippers. The dexterous hands can perform human-like manipulations such as twisting and screwing, and autonomously select actions based on tactile feedback [36]. Although grippers can also be equipped with tactile sensors, they can only detect basic grasp status rather than enabling active environmental interaction. By contrast, dexterous hands can use rich tactile sensing to compensate for limited visual perception (e.g., blocked cameras) and interact adaptively with the environment. Thus, only dexterous hands are capable of addressing all challenges in robotic manipulation.
@@
However, learning manipulation skills for multi-fingered hands remains significantly more difficult than for simple two-fingered grippers. Compared with the simple two-fingered gripper, the learning of manipulation skills for multi-fingered dexterous hands is particularly difficult. First, the multi-fingered dexterous hand has a higher degree of freedom (DoF) and a more variable interaction space, which greatly increases the difficulty of searching in high-dimensional space for algorithms such as RL. Second, in the process of performing manipulation tasks, the multi-fingered dexterous hand has multiple contact points with objects, leading to more complex and diverse mechanical properties. This further makes the learning of skills for multi-fingered dexterous hands more difficult.
@@ IV The Data Collection Paradigm for Dexterous Manipulation
ActionSense [46] Multi-Fingered Hand An expanding dataset focused on daily tasks in a kitchen environment featuring wearable sensors for body and finger tracking, forearm muscle activity, tactile information, and eye tracking with first-person video.
@@ IV-A Data Generation Based on Simulation Platform
Faced with complex, diverse, and highly uncertain application scenarios, traditional learning methods that rely on "robot-environment" interaction to obtain intelligence often require millions of iterations to learn useful skills, which is generally inefficient. Therefore, the most common approach is to generate a large amount of data through simulation platforms to improve efficiency. Data generation based on simulation platforms has some unique advantages. First, generating data on the simulation platform can avoid the high cost and potential danger of repeated experiments in the real world. Second, the simulation platform can efficiently, cost-effectively and repeatably generate a large amount of pedigree data with comparative significance. This can enhance the diversity of datasets and improve the quality of generated data. Finally, the simulation platform can accelerate the speed of data collection that may take longer in the real physical world by changing the flow rate of time in simulation environment.
At present, there have developed some simulation datasets based on physics engines such as Genesis[52], Isaac Sim, PyBullet[53], and MoJoCo[54]. For example, the GraspM3[38] simulation dataset that provides millions of grasping trajectories covering 8,152 objects and a generative robot agent RoboGen[39] that can generate data infinitely. Wang et al.[40] generated a large-scale simulation dataset DexGraspNet for Shadow Hand, which contains 1.32 million grasping data of 5,355 objects in 133 categories. Shanghai Artificial Intelligence Laboratory[41] released a city-level embodied intelligent simulation platform GRPtopia, which can provide 89 functional scenarios and 100,000 high-quality interactive data for robot training.
Meanwhile, the tactile feedback serving as the crucial information for dexterous manipulation has also attracted increasing attention in data generation and collection. Wang et al. [55] designed a simulator, named TACTO, for vision-based tactile sensors. This simulator allows to render realistic high-resolution touch readings in dexterous manipulation at hundreds of frames per second, and can be easily configured to simulate different vision-based tactile sensors. Zhong et al. [56] proposed a TactGen for cross-modal visuo-tactile data generation, which can achieve zero-shot sim-to-real transfer without any real-world data for domain adaptation. Hong et al. [57] collected a large-scale multisensory interaction dataset, named Multisensory Universe, comprising 500k data by deploying an LLM-powered embodied agent to engage with the 3D environment. Despite progress in generating tactile data, great potential remains for creating a dexterous manipulation simulation dataset with tactile sensing.
@@ IV-B Data Collection from Human Demonstration
The ultimate goal of robotic manipulation is to endow robots with human-like dexterous manipulation capabilities. Therefore, learning manipulation skills directly from human demonstration actions has become an attractive solution. Although the scale of data collection based on human motion demonstrations cannot compare with the data generated by simulation, it significantly reduces the difficulty compared to repeated trials with real robots, and can expand the scope of data collection without requiring robot hardware. At the same time, compared with the data generated by simulation, the human demonstration data is real interaction data produced in the physical environment, which can greatly reduce the Sim2Real gap.
For instance, Bahl et al.[42][43] proposed to extract human motions and key interaction elements from billions of video data that are already widely available in Internet to achieve manipulation data collection based on human demonstration. Liu et al.[44][45] also developed a vision-tactile fusion human motion capture system, constructing the VTDexManip dataset. DelPreto et al. [46] built a wearable data collection framework for recording synchronized multimodal data through body trackers, first-person video with attention estimates, muscle activity sensors, and custom tactile gloves. They used this system to collect an expanding dataset focused on daily tasks in a kitchen environment. Zhao et al. [58] used a wearable tool-like parallel hand exoskeleton with visual-based tactile sensors to collect data from human demonstration by grasping 15 different daily objects. These works enhance robots’ manipulation capabilities in real physical environments.
@@ V The Learning Framework for Dexterous Manipulation Skills
BC / — — [71] Combine behavior classification and geometric augmentation to enable effective behavioral cloning from mixed-quality demonstration datasets.
BC / DIME [72] Capture demonstrations from the real-time visual stream of a human operator’s hand, utilize these demonstrations to teleoperate a robotic hand, and learn IL policies based on the robotic hand’s actions.
BC / DIH-Tele [73] Utilized proprioception, demonstration action sequence, and tactile information to train IL policies.
BC / ViTacFormer [74] Utilized proprioception, demonstration action sequence, visual, and tactile information to train IL policies.
BC / DexForce [75] Utilize force-informed actions to train IL policies. These force-informed actions are computed based on the contact forces which measured during kinesthetic demonstrations.
BC / DexSkills [36] Utilize tactile features acquired from the demonstrations to predict the on going primitives of the long-horizontal dexterous manipulation tasks.
IRL [76] Learn the reward function using samples obtained from demonstrations of desired behaviors. Use statistical tools for random sample generation and reward normalization to reduce demonstration-dependent reward bias.
GAIL [77] An adversarial imitation learning algorithm where a policy learns to mimic expert behavior by tricking a discriminator which tries to distinguish it from the expert.
AIRL [76] Recover policies and interpretable reward functions simultaneously within an adversarial learning framework. thereby providing strong transferability and generalization capability.
@@
— — [82] Utilize visual tactile information to train RL policies. Policies that obtained by RL can be generalized to unseen objects.
@@
— — [85] Leverage high-dimensional tactile information from interactions to train a DRL policy to control the pose of a thin cylindrical object in a tracking task.
DexTouch [86] Rely solely on tactile sensing to perform manipulation tasks, with vision excluded.
VTAO-BiManip [87] leverage visual–tactile–action pretraining with object understanding to facilitate reinforcement learning and achieve human-like bi-manual manipulation.
@@
DAPG [89] DRL is augmented with a small number of human demonstrations.
DexH2R [90] Combine human hand motion re-targeting with a task-oriented residual action policy, improve task performance by bridging the embodiment gap between human and robotic dexterous hands.
@@
LEGION [92] A Bayesian-inspired lifelong RL framework integrates language embeddings to enable robots to continuously accumulate and reuse knowledge from sequential tasks, advancing generalizable robotic intelligence.
@@ V-A Imitation Learning
The second subcategory of IL employs deep learning to train policy networks that mimic human expert decisions. Unlike RL, this approach does not rely on environmental interaction or rewards but learns from pre-collected datasets. Common methods include Behavior Cloning (BC)[71, 72, 73, 74, 75, 36], Inverse reinforcement learning (IRL) [93, 76], Adversarial Inverse Reinforcement Learning (AIRL) [76], and Generative Adversarial Imitation Learning (GAIL) [76, 77]. Among these three methods, BC has generally demonstrated superior performance [71] and has therefore received considerable attention in recent years. In [72], demonstrations for a nearest neighbor-based IL is generated by a teleoperation system. In this teleoperation system, a vision system is used to extract the movements of the operator’s fingertips. These movements are then used to control the movements of a robotic hand. The method proposed in [72] demonstrates excellent performance in complex in-hand manipulation tasks. However, for these contact-rich tasks, no tactile information is incorporated. Incorporating tactile feedback could potentially further improve the learning outcomes. In [73], tactile images along with the positional information are used as input to generate action sequences. In [74]. except for the positional information and the tactile information, visual information are also used. The multi-modal representation enables IL for multi-fingered dexterous hands, enabling precise and adaptive manipulation. In [75], the IL policies are trained on force-informed actions. These force-informed actions are computed based on the contact forces which measured during kinesthetic demonstrations. In [36], tactile features acquired from the demonstrations are used to predict the on going primitives of the long-horizontal dexterous manipulation tasks. A key limitation is that these methods only learn static behaviors from the demonstration data and cannot surpass human performance as RL might. Additionally, their performance degrades significantly when encountering states outside the training distribution, highlighting robustness issues.
@@ V-B Reinforcement Learning
RL has become the dominant approach for robotic dexterous manipulation skill learning, enabling agents to interact with their environment and refine policies through reward feedback. Recent researches have developed specialized simulation environments (e.g., Bi-DexHands for bi-manual manipulation)[78], and RL methods for robotic control. Compared with traditional feedback control methods, RL has successfully tackled challenging manipulation skills[79, 94, 80]. However, applying RL to multi-fingered dexterous hands presents unique challenges due to the high degrees of freedom and complex contact dynamics[95]. Pure RL that only relies on agent-environment often suffers from sparse rewards and low sample efficiency, typically requiring millions of iterations to learn useful behaviors. The exploration process can also lead to incoherent or unsafe actions, making it difficult to achieve desired learning outcomes. To enhance learning efficiency, recent studies have utilized features extracted by pre-trained models to guide downstream robot tasks[96, 81]. The use of existing pre-trained models, combined with vast internet data, enables rapid and efficient learning of new tasks. In robotic RL, this pre-trained paradigm offers a key advantage: robots can acquire new skills with minimal human intervention by leveraging publicly available data and pre-trained models.
Same as IL, tactile information can be quite useful. Su et al. employed RL to train a robot to pivot an object to a target orientation using only proprioceptive and tactile information[82]. Liu et al. proposed a novel framework called TactileAIRL for skill learning in robotic dexterous manipulation. Model-based techniques and intrinsic curiosity are integrated into the RL process, the method leverages vision-based tactile sensing to extract meaningful contact features. This design makes TactileAIRL scalable to many manipulation tasks learning involving tactile feedback[83]. Zhou et al. presented an RL framework called T-TD3 for stable grasping of deformable objects using tactile prior information[84]. Aslam et al. presented a robot called DartBot that integrates tactile exploration and RL to achieve robust throwing skills in object transfer through throwing task[88]. Hu et al. [85] proposed leveraging high-dimensional tactile information from interactions between a thin stick and the fingertips of a three-fingered robotic hand to train a DRL policy for a tracking task. The policies were trained in simulation and successfully transferred to real-world experiments. Lee et al. [86] proposed a high-DoF dexterous robotic system with tactile-based blind manipulation capabilities. The system relies solely on tactile sensing to perform manipulation tasks, with vision excluded. Sun et al. [87] proposed leveraging visual–tactile–action pretraining with object understanding to facilitate reinforcement learning and achieve human-like bi-manual manipulation.